OpenStudy (anonymous):

would the series comparison for Sigma notation going from 2 to infinity of 1/n(n^2 - 1)^1/2 be equal to 1/n^2?

OpenStudy (anonymous):

or how am I able to find if the equation converges or diverges?

OpenStudy (anonymous):

it converges I think !

OpenStudy (anonymous):

you know about Riemmann's series ?!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I tried using series comparison but it didn't work cause i don't know if I'm comparing it with the right one

OpenStudy (anonymous):

okk compare it with \[1/n^(3/2) \]

OpenStudy (anonymous):

your serie is \[\sum_{}^{} 1/\sqrt{n(n^2-1)}\] ???!!

OpenStudy (anonymous):

no, my serie is \[\sum_{2}^{infinity} 1/x \sqrt{x^2 -1}\]

OpenStudy (anonymous):

yeaaah ! compare it wi 1/x^2 \[1/x(\sqrt{x^2-1})\approx( infinity)1/x^2\] \[\lim_{x \rightarrow infinity}1/x \sqrt{x^2-1}/(1/x^2)=1\]

OpenStudy (anonymous):

I tried it with the integral test, and after a while of playing around I managed to get: \[ \int_2^\infty =\frac{1}{x\sqrt{x^2-1}} = \tan^{-1}(\sqrt{x^2-1}) \] which converges for large values of x.

OpenStudy (anonymous):

i compared it with that same one and by means of p-series it converges because p=2 which \[2>1\] and then proceeded to prove it by inequalities but I'm confused with the result \[(x^2<0)\]

OpenStudy (anonymous):

yeaaah ; I see...but inequalities...how ?

OpenStudy (anonymous):

based on the formulaf you do it by the integral test its not tan−1 instead its sec−1

OpenStudy (anonymous):

because you have to prove that if the one you compare it with converges than the original also converges

OpenStudy (anonymous):

yeaaah...1/x^2 you said that it converges .?! you have a problem with that ?

OpenStudy (anonymous):

Moni I understand your point, but you can derive my formula and it should work.

OpenStudy (anonymous):

or maybe I am really missing an important step, I won't exclude that.

OpenStudy (anonymous):

\[ f(x) = \tan^{-1}(\sqrt{x^2-1}) \\ f'(x) = \frac{1}{x^2} \cdot \frac{1}{2\sqrt{x^2-1}} \cdot 2x = \frac{1}{x\sqrt{x^2-1}} \]

OpenStudy (anonymous):

its ok don't worry about it guys! i managed to find it THANKS so much!!! i diet click the points together....

OpenStudy (anonymous):

w8 , tell me how you diid it ?! please ! :)

OpenStudy (anonymous):

by doing the inequalities \[1/x \sqrt{x^2-1}< 1/x^2\]

OpenStudy (anonymous):

sorry but it's false ! :) it's true only if x is a negatif number !

OpenStudy (anonymous):

then you work for each side making it \[x^2< x \sqrt{x^2-1}\] and than you square each side to eliminate the square root... and you have \[x^4< x^2(x(2-1)\] you multiply and then cancel the x^4 and the result will be \[0<-x^2\]

OpenStudy (anonymous):

the second part should be \[x^4<x^2(x^2-1)\]

OpenStudy (anonymous):

no ! \[x^2-1\le x^2\] then \[\sqrt{x^2-1} \le \sqrt{x^2}=x \] you multiply by x for both sides because x is a positive number . am I right ?!

OpenStudy (anonymous):

if you want to do it with inequalities think about something else...

OpenStudy (anonymous):

let me keep trying and if not i'll com back... thanks though

OpenStudy (anonymous):

you are a real fighter ! I'm thinking about it too...:D

OpenStudy (anonymous):

@monibea you know what does"is asymptotically equivalent to" mean ?! because with inequalities...things gets more complicated !